OpenStudy (anonymous):

Can someone help me to solve this? square root of 2x^3y^3 times cube root of 4xy^2 Thank You!

OpenStudy (australopithecus):

Use the following rules \[(xy)^{1/2} = x^{1/2}y^{1/2}\] \[(x^2)^{1/2} = (x)^{2\frac{1}{2}} = (x)^{\frac{2}{1}\frac{1}{2}} = (x)^{\frac{(2)1}{2}}\]

OpenStudy (australopithecus):

where \[(x)^{1/2} = \sqrt{x}\] and \[(x)^{1/3} = \sqrt[3]{x}\]

OpenStudy (anonymous):

but the answer says: 2xy^2 6th root of 2x^5y

OpenStudy (australopithecus):

just to clarify in a case with a constant \[(2x^2)^{1/2} = 2^{1/2}(x^2)^{1/2}\]

OpenStudy (australopithecus):

I'm showing yo how to simplify radicals

OpenStudy (anonymous):

i know how to simplify, but my answer doesn't match with the book

OpenStudy (australopithecus):

what is your answer?

OpenStudy (anonymous):

the answer is: 2xy^2 6th root of 2x^5y

OpenStudy (anonymous):